the average salary for registered nurses is $3500 less than three times the average salary for janitors. combined, their average salaries are $74,060. Determine the average salary for each of these jobs.
Please help i am really bad at word problems please show me how you got the anwser
thank you. holly b

Answer by richwmiller(17219)

n=3j-3500
n+j=74060
n=74060-j
n=3j-3500
set these two equal since they are both equal to n
74060-j=3j-3500
add j to both sides
add 3500 to both sides
77560=4j
j = 19390
n=74060-j
n = 54670